# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 11 Acknowledgments
The author would like to thank Udai Parvathaneni, Nathan Sheng, JD Maturen, Kamil Yusubov, and Luke Sterle from Gauntlet for helpful dicussions around how to quantify realistic ADL scenarios (such as October 10, 2025). The author would also like to thank Matheus V. X. Ferreira, Guillermo Angeris, Victor Xu, and Vinayak Kurup for helpful discussions. Finally, the author really appreciates the data provided by SonarX, Hydromancer, and Mauricio Trujillo.
Notation and conventions
@@
  • Shocks are indexed by $t=1,2,\dots,T$; winners $W_{t}$ with equities $e_{t,i}>0$ and losers $L_{t}$.
  • Deficit $D_{t}\geq 0$; severity $\theta_{t}\geq 0$; haircut vector $h_{t}\in[0,1]^{|W_{t}|}$ with survivors $s_{t,i}=(1-h_{t,i})e_{t,i}$.
  • Per–round budget $B_{t}=\min\{\theta_{t}D_{t},\ \sum_{i}w_{t,i}\}$ for severity policies and $B_{t}=w_{t}^{\top}h_{t}$ for vector policies; $B_{t}^{\star}=\min\{D_{t},\sum_{i}w_{t,i}\}$.
@@
Appendix A Liquidations, Autodeleveraging, and Insurance Funds
A.1 Bankruptcy Price Example
@@
A.2 Liquidation Price Example
@@
A.3 Execution Price Example
@@
A.4 Liquidation Costs Example
@@
  • Binance: $(\tau^{fix},\phi^{mark},\phi^{exec})=(0,\,40\,\mathrm{bps},\,0)$ [BinanceFuturesInsuranceFund]. Then $\tau=0.0040\cdot 1.30\cdot 0.50=0.0026$.
  • Hyperliquid: $(\tau^{fix},\phi^{mark},\phi^{exec})=(0,\,20\,\mathrm{bps},\,10\,\mathrm{bps})$ [HyperliquidDocsLiquidations]. Then $\tau=0.0020\cdot 1.30\cdot 0.50+0.0010\cdot 1.32\cdot 0.50\approx 0.00130+0.00066=0.00196$.
@@
A.5 Liquidation Strategy Example
@@ A.7 Anatomy of a Liquidation
A.8 Optimal Capital Structure Derivation
@@
Appendix B Moral Hazard and Extreme Value Analysis
@@
B.1 Setup and Assumptions
We work in the large-market limit ($n\to\infty$) under the heavy-tailed assumptions characteristic of crypto markets. Recall that $D^{\pi}_{T}=\theta_{\pi}D_{T}$ is the total socialized loss and $\Delta^{\pi}_{T}=\theta_{\pi}\Delta_{T}$ is the maximum socialized shortfall under policy $\pi$. The survivor of the top winner is denoted $\omega^{\pi}_{T}$.
@@ B.1 Setup and Assumptions
The right tail of the winner equity distribution $\bar{F}_{+}(x)$ and the right tail of the loser shortfall distribution $\bar{F}_{-}(x)$ are regularly varying with indices $\alpha_{+}>0$ and $\alpha_{-}>0$, respectively. That is, $\bar{F}_{\pm}(x)=L_{\pm}(x)x^{-\alpha_{\pm}}$ where $L_{\pm}$ are slowly varying functions.
@@
  • Aggregates: $W_{T}/n\xrightarrow{p}\mu_{+}$ and $D_{T}/n\xrightarrow{p}\mu_{-}$, for constants $\mu_{\pm}\in(0,\infty)$.
  • Extremes: The maximum winner $\omega_{T}$ scales as $b_{k_{n}}^{+}=F_{+}^{\leftarrow}(1-1/n)$, and the maximum loser shortfall $\Delta_{T}$ scales as $b_{m_{n}}^{-}=F_{-}^{\leftarrow}(1-1/n)$.
B.2 Optimal Control of Moral Hazard
B.3 Queue maximizes top-winner damage
@@
Proposition B.1 (Queue Minimizes Top Survivor).
Fix a budget $H=D^{\pi}_{T}$. Let $\omega_{T}$ be the equity of the largest winner. For any feasible haircut vector $h$ satisfying $\sum h_{i}e_{i}=H$, the top-winner survivor $\omega^{\pi}_{T}$ satisfies
$\omega^{\pi}_{T}\;\geq\;\max\{\omega_{T}-H,0\}.$
Equality is attained by the Queue rule, which sets the haircut on the top winner to $h_{(1)}=\min(H/\omega_{T},1)$ and others to 0 (until $h_{(1)}$ saturates).
@@
Let $h_{(1)}$ be the haircut applied to the top winner. Since $h_{i}e_{i}\geq 0$ for all $i$, we have $h_{(1)}\omega_{T}\leq\sum_{i}h_{i}e_{i}=H$. The survivor is $\omega^{\pi}_{T}=\omega_{T}-h_{(1)}\omega_{T}\geq\omega_{T}-H$. Since equity cannot be negative, $\omega^{\pi}_{T}\geq\max\{\omega_{T}-H,0\}$. The Queue rule greedily allocates the budget to the largest position, achieving $h_{(1)}\omega_{T}=\min(H,\omega_{T})$, thus attaining the lower bound. ∎
Corollary B.2 (Minimality of PTSR/PMR).
@@
The Queue rule minimizes moral hazard but concentrates the entire loss on the top winner (extreme inequality). In contrast, the Pro-Rata rule spreads the loss proportionally across all winners, prioritizing smoothness (treating similar positions similarly) over minimizing the top survivor’s burden. For $H\leq\omega_{T}$, the survivor gap is
$\omega^{\mathrm{PR}}_{T}-\omega^{\mathrm{Queue}}_{T}\;=\;H\left(1-\frac{\omega_{T}}{W_{T}}\right).$
This gap scales linearly with the budget $H$, quantifying the “cost of fairness”: by choosing the smoother Pro-Rata allocation, the system allows the top winner to retain more profit than is strictly necessary to cover the deficit.
B.4 Asymptotic Scaling Results
@@
Theorem B.3 (PTSR scaling).
Under Assumptions A and B, for any policy $\pi$ with severity $\theta_{n}$ where $\omega^{\pi}_{T}\sim\omega_{T}$ (e.g., Pro-Rata with $H\ll W_{T}$), the PTSR scales as
$\mathsf{PTSR}_{T}(\pi)\;\asymp_{p}\;\frac{b_{k_{n}}^{+}}{\theta_{n}n}.$
@@
By definition, $\mathsf{PTSR}_{T}(\pi)=\mathop{\bf E{}}[\omega^{\pi}_{T}/D^{\pi}_{T}]$. Under the hypothesis, the numerator scales as $\omega_{T}\sim b_{k_{n}}^{+}$. The denominator is $D^{\pi}_{T}=\theta_{n}D_{T}$. By the LLN, $D_{T}\sim\mu_{-}n$, so $D^{\pi}_{T}\sim\theta_{n}\mu_{-}n$. Thus, the ratio scales as $b_{k_{n}}^{+}/(\theta_{n}n)$. Using bounded convergence for the expectation yields the result. ∎
@@
The behavior of PTSR depends critically on the tail class of winner equities:
@@
  • Exponential/Gaussian (Light) Tails: $b_{k_{n}}^{+}\asymp(\log n)^{\gamma}$. Here $\mathsf{PTSR}_{T}\asymp(\log n)^{\gamma}/(n\theta_{n})$. Since polylog growth is slower than linear, moral hazard vanishes rapidly for any non-vanishing severity $\theta_{n}$, as the aggregate deficit overwhelms the largest individual winner.
Theorem B.4 (PMR Scaling).
Assume winner equities have mass $\ell_{n}^{+}$ and loser deficits have mass $\ell_{n}^{-}$ (representing total leverage), and that the underlying normalized distributions satisfy Assumption A. The PMR scales as:
$\mathsf{PMR}_{T}(\pi)\;\asymp_{p}\;\frac{1}{\theta_{n}}\cdot\frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot\frac{b_{k_{n}}^{+}}{b_{m_{n}}^{-}}\;\asymp\;\frac{1}{\theta_{n}}\cdot\frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot n^{\frac{1}{\alpha_{+}}-\frac{1}{\alpha_{-}}}.$
@@
We have $\mathsf{PMR}_{T}(\pi)=\mathop{\bf E{}}[\omega^{\pi}_{T}/\Delta^{\pi}_{T}]$. The top winner scales with total winner leverage mass: $\omega_{T}\sim\ell_{n}^{+}b_{k_{n}}^{+}$. The maximum loser shortfall scales with total loser leverage mass: $\Delta_{T}\sim\ell_{n}^{-}b_{m_{n}}^{-}$. The budget balance condition implies $\Delta^{\pi}_{T}=\theta_{n}\Delta_{T}$. Thus, the ratio scales as
@@
Theorem B.4 decomposes moral hazard into three components: (1) Policy Severity ($1/\theta_{n}$): Lower severity amplifies PMR. (2) Leverage Imbalance ($\ell_{n}^{+}/\ell_{n}^{-}$): If the winning side holds more leverage mass, PMR increases. (3) Tail Risk ($n^{1/\alpha_{+}-1/\alpha_{-}}$): Heavier winner tails relative to losers drive PMR divergence. This decomposition highlights that even with fair tails ($\alpha_{+}=\alpha_{-}$), a systemic leverage imbalance ($\ell_{n}^{+}\gg\ell_{n}^{-}$) can sustain a high PMR. Specifically, if the exchange allows winners to be significantly more leveraged than losers (a "risk-on" imbalance), the top winner’s survival will systematically outstrip the worst-case socialized loss, creating a persistent moral hazard where maximal profits are privatized while maximal losses are capped.
B.5 Relationship to Classical Risk Measures
@@
Proposition B.5 (Queue minimizes VaR/ES of the top survivor).
@@
The pointwise lower bound $\omega^{\pi}_{T}\geq(\omega_{T}-h)_{+}$ follows from the budget constraint and nonnegativity of haircuts, as in Proposition B.1. Monotonicity of risk measures implies that if $X\geq Y$ almost surely, then $\mathrm{VaR}_{\alpha}(X)\geq\mathrm{VaR}_{\alpha}(Y)$ and $\mathrm{ES}_{\alpha}(X)\geq\mathrm{ES}_{\alpha}(Y)$. For the identities, observe that $x\mapsto(x-h)_{+}$ is nondecreasing; hence quantiles shift: $\mathrm{VaR}_{\alpha}((X-h)_{+})=\max\{\mathrm{VaR}_{\alpha}(X)-h,0\}$. The ES identity follows from the Kusuoka representation [Kusuoka2001] $\mathrm{ES}_{\alpha}(Z)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\mathrm{VaR}_{u}(Z)\,du$ applied to $Z=(X-h)_{+}$. ∎
@@
For a random budget $H=\theta_{n}D_{T}$, apply Theorem B.5 conditionally on $H$ to conclude that Queue minimizes the conditional VaR/ES of the top survivor at every tail level. When $\alpha_{+}>1$ (finite mean winners), the tail-equivalence property of regularly varying distributions yields
@@
B.6 Randomized constructions for moral-hazard examples
Extreme-value moral-hazard (principal–agent).

### v3

@@ 11 Acknowledgments
The author would like to thank Udai Parvathaneni, Nathan Sheng, JD Maturen, Kamil Yusubov, and Luke Sterle from Gauntlet for helpful discussions around how to quantify realistic ADL scenarios (such as October 10, 2025). The author would also like to thank Nagu Thogiti, Victor Xu, Yi Sun, Anirudh Pai, Kshitij Kulkarni, Theo Diamandis, Matheus V. X. Ferreira, Guillermo Angeris, and Vinayak Kurup for helpful discussions. Finally, the author really appreciates the help with procuring and cleaning Hyperliquid data provided by SonarX, Hydromancer, and Mauricio Trujillo (@conejocapital).
Appendix
Appendix A Notation and conventions
@@
  • Shocks are indexed by $t=1,2,\dots,T$; winners $W_{t}$ with endowments $w_{t,i}>0$ (under PNL-only, $w_{t,i}=(\mathrm{PNL}_{t,i})_{+}$) and losers $L_{t}$ with negative equity.
  • Deficit $D_{t}\geq 0$ (computed from loser negative equity); severity $\theta_{t}\geq 0$; haircut vector $h_{t}\in[0,1]^{|W_{t}|}$ with endowment survivors $s_{t,i}=(1-h_{t,i})w_{t,i}$.
  • Per–round budget $B_{t}=\min\{\theta_{t}D_{t},\ \sum_{i}w_{t,i}\}$ for severity policies and $B_{t}=w_{t}^{\top}h_{t}$ for vector policies; $B_{t}^{\star}=\min\{D_{t},\sum_{i}w_{t,i}\}$ (budget constraint in endowment space).
@@
Appendix B Liquidations, Autodeleveraging, and Insurance Funds
B.1 Bankruptcy Price Example
@@
B.2 Liquidation Price Example
@@
B.3 Execution Price Example
@@
B.4 Liquidation Costs Example
@@
  • Binance: $(\tau^{fix},\phi^{mark},\phi^{exec})=(0,\,40\,\mathrm{bps},\,0)$ [22]. Then $\tau=0.0040\cdot 1.30\cdot 0.50=0.0026$.
  • Hyperliquid: $(\tau^{fix},\phi^{mark},\phi^{exec})=(0,\,20\,\mathrm{bps},\,10\,\mathrm{bps})$ [69]. Then $\tau=0.0020\cdot 1.30\cdot 0.50+0.0010\cdot 1.32\cdot 0.50\approx 0.00130+0.00066=0.00196$.
@@
B.5 Liquidation Strategy Example
@@ B.7 Anatomy of a Liquidation
B.8 Optimal Capital Structure Derivation
@@
Appendix C Moral Hazard and Extreme Value Analysis
@@
C.1 Setup and Assumptions
We work in the large-market limit ($n\to\infty$) under the heavy-tailed assumptions characteristic of crypto markets. Recall that $D^{\pi}_{T}=\theta_{\pi}D_{T}$ is the total socialized loss and $\Delta^{\pi}_{T}=\theta_{\pi}\Delta_{T}$ is the maximum socialized shortfall under policy $\pi$. The survivor of the top winner’s endowment is denoted $\upsilon^{\pi}_{T}=\max_{i}(w_{i,T}-x_{\pi,i})_{+}$ (under PNL-only, this is the maximum post-ADL positive PNL).
@@ C.1 Setup and Assumptions
The right tail of the winner endowment distribution $\bar{F}_{+}(x)$ (under PNL-only, the distribution of positive PNL) and the right tail of the loser shortfall distribution $\bar{F}_{-}(x)$ are regularly varying with indices $\alpha_{+}>0$ and $\alpha_{-}>0$, respectively. That is, $\bar{F}_{\pm}(x)=L_{\pm}(x)x^{-\alpha_{\pm}}$ where $L_{\pm}$ are slowly varying functions.
@@
  • Aggregates: $U_{T}/n\xrightarrow{p}\mu_{+}$ (total endowment capacity) and $D_{T}/n\xrightarrow{p}\mu_{-}$ (total deficit), for constants $\mu_{\pm}\in(0,\infty)$.
  • Extremes: The maximum winner endowment $\upsilon_{T}$ scales as $b_{k_{n}}^{+}=F_{+}^{\leftarrow}(1-1/n)$, and the maximum loser shortfall $\Delta_{T}$ scales as $b_{m_{n}}^{-}=F_{-}^{\leftarrow}(1-1/n)$.
C.2 Optimal Control of Moral Hazard
C.3 Queue maximizes top-winner damage
@@
Proposition C.1 (Queue Minimizes Top Survivor).
Fix a budget $H=D^{\pi}_{T}$. Let $\upsilon_{T}$ be the endowment of the largest winner (under PNL-only, the maximum positive PNL). For any feasible haircut vector $h$ satisfying $\sum_{i}h_{i}w_{i,T}=H$ (budget constraint in endowment space), the top-winner endowment survivor $\upsilon^{\pi}_{T}$ satisfies
$\upsilon^{\pi}_{T}\;\geq\;\max\{\upsilon_{T}-H,0\}.$
Equality is attained by the Queue rule, which sets the haircut on the top winner to $h_{(1)}=\min(H/\upsilon_{T},1)$ and others to 0 (until $h_{(1)}$ saturates).
@@
Let $h_{(1)}$ be the haircut applied to the top winner’s endowment. Since $h_{i}w_{i,T}\geq 0$ for all $i$, we have $h_{(1)}\upsilon_{T}\leq\sum_{i}h_{i}w_{i,T}=H$. The survivor is $\upsilon^{\pi}_{T}=\upsilon_{T}-h_{(1)}\upsilon_{T}\geq\upsilon_{T}-H$. Since endowment cannot be negative, $\upsilon^{\pi}_{T}\geq\max\{\upsilon_{T}-H,0\}$. The Queue rule greedily allocates the budget to the largest endowment position, achieving $h_{(1)}\upsilon_{T}=\min(H,\upsilon_{T})$, thus attaining the lower bound. ∎
Corollary C.2 (Minimality of PTSR/PMR).
@@
The Queue rule minimizes moral hazard but concentrates the entire loss on the top winner (extreme inequality). In contrast, the Pro-Rata rule spreads the loss proportionally across all winners, prioritizing smoothness (treating similar positions similarly) over minimizing the top survivor’s burden. For $H\leq\upsilon_{T}$, the survivor gap is
$\upsilon^{\mathrm{PR}}_{T}-\upsilon^{\mathrm{Queue}}_{T}\;=\;H\left(1-\frac{\upsilon_{T}}{U_{T}}\right),$
where $U_{T}=\sum_{i}w_{i,T}$ is the total endowment capacity. This gap scales linearly with the budget $H$, quantifying the “cost of fairness”: by choosing the smoother Pro-Rata allocation, the system allows the top winner to retain more profit capacity than is strictly necessary to cover the deficit.
C.4 Asymptotic Scaling Results
@@
Theorem C.3 (PTSR scaling).
Under Assumptions A and B, for any policy $\pi$ with severity $\theta_{n}$ where $\upsilon^{\pi}_{T}\sim\upsilon_{T}$ (e.g., Pro-Rata with $H\ll U_{T}$), the PTSR scales as
$\mathsf{PTSR}_{T}(\pi)\;\asymp_{p}\;\frac{b_{k_{n}}^{+}}{\theta_{n}n},$
where $\upsilon^{\pi}_{T}$ is the maximum post-ADL endowment.
@@
By definition, $\mathsf{PTSR}_{T}(\pi)=\mathop{\bf E{}}[\upsilon^{\pi}_{T}/D^{\pi}_{T}]$. Under the hypothesis, the numerator scales as $\upsilon_{T}\sim b_{k_{n}}^{+}$ (the maximum endowment, which under PNL-only is the maximum positive PNL). The denominator is $D^{\pi}_{T}=\theta_{n}D_{T}$. By the LLN, $D_{T}\sim\mu_{-}n$, so $D^{\pi}_{T}\sim\theta_{n}\mu_{-}n$. Thus, the ratio scales as $b_{k_{n}}^{+}/(\theta_{n}n)$. Using bounded convergence for the expectation yields the result. ∎
@@
The behavior of PTSR depends critically on the tail class of winner endowments (under PNL-only, positive PNL):
@@
  • Exponential/Gaussian (Light) Tails: $b_{k_{n}}^{+}\asymp(\log n)^{\gamma}$. Here $\mathsf{PTSR}_{T}\asymp(\log n)^{\gamma}/(n\theta_{n})$. Since polylog growth is slower than linear, moral hazard vanishes rapidly for any non-vanishing severity $\theta_{n}$, as the aggregate deficit overwhelms the largest individual winner’s endowment.
Theorem C.4 (PMR Scaling).
Assume winner endowments scale with leverage mass $\ell_{n}^{+}$ and loser deficits scale with leverage mass $\ell_{n}^{-}$ (representing total leverage), and that the underlying normalized distributions satisfy Assumption A. The PMR scales as:
$\mathsf{PMR}_{T}(\pi)\;\asymp_{p}\;\frac{1}{\theta_{n}}\cdot\frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot\frac{b_{k_{n}}^{+}}{b_{m_{n}}^{-}}\;\asymp\;\frac{1}{\theta_{n}}\cdot\frac{\ell_{n}^{+}}{\ell_{n}^{-}}\cdot n^{\frac{1}{\alpha_{+}}-\frac{1}{\alpha_{-}}},$
where $\upsilon^{\pi}_{T}$ is the maximum post-ADL endowment.
@@
We have $\mathsf{PMR}_{T}(\pi)=\mathop{\bf E{}}[\upsilon^{\pi}_{T}/\Delta^{\pi}_{T}]$. The top winner’s endowment scales with total winner leverage mass: $\upsilon_{T}\sim\ell_{n}^{+}b_{k_{n}}^{+}$ (under PNL-only, this is the maximum positive PNL). The maximum loser shortfall scales with total loser leverage mass: $\Delta_{T}\sim\ell_{n}^{-}b_{m_{n}}^{-}$. The budget balance condition implies $\Delta^{\pi}_{T}=\theta_{n}\Delta_{T}$. Thus, the ratio scales as
@@
Theorem C.4 decomposes moral hazard into three components: (1) Policy Severity ($1/\theta_{n}$): Lower severity amplifies PMR. (2) Leverage Imbalance ($\ell_{n}^{+}/\ell_{n}^{-}$): If the winning side holds more leverage mass, PMR increases. (3) Tail Risk ($n^{1/\alpha_{+}-1/\alpha_{-}}$): Heavier winner tails relative to losers drive PMR divergence. This decomposition highlights that even with fair tails ($\alpha_{+}=\alpha_{-}$), a systemic leverage imbalance ($\ell_{n}^{+}\gg\ell_{n}^{-}$) can sustain a high PMR. Specifically, if the exchange allows winners to be significantly more leveraged than losers (a "risk-on" imbalance), the top winner’s survival will systematically outstrip the worst-case socialized loss, creating a persistent moral hazard where maximal profits are privatized while maximal losses are capped.
C.5 Relationship to Classical Risk Measures
@@
Proposition C.5 (Queue minimizes VaR/ES of the top survivor).
@@
The pointwise lower bound $\omega^{\pi}_{T}\geq(\omega_{T}-h)_{+}$ follows from the budget constraint and nonnegativity of haircuts, as in Proposition C.1. Monotonicity of risk measures implies that if $X\geq Y$ almost surely, then $\mathrm{VaR}_{\alpha}(X)\geq\mathrm{VaR}_{\alpha}(Y)$ and $\mathrm{ES}_{\alpha}(X)\geq\mathrm{ES}_{\alpha}(Y)$. For the identities, observe that $x\mapsto(x-h)_{+}$ is nondecreasing; hence quantiles shift: $\mathrm{VaR}_{\alpha}((X-h)_{+})=\max\{\mathrm{VaR}_{\alpha}(X)-h,0\}$. The ES identity follows from the Kusuoka representation [78] $\mathrm{ES}_{\alpha}(Z)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\mathrm{VaR}_{u}(Z)\,du$ applied to $Z=(X-h)_{+}$. ∎
@@
For a random budget $H=\theta_{n}D_{T}$, apply Theorem C.5 conditionally on $H$ to conclude that Queue minimizes the conditional VaR/ES of the top survivor at every tail level. When $\alpha_{+}>1$ (finite mean winners), the tail-equivalence property of regularly varying distributions yields
@@
C.6 Randomized constructions for moral-hazard examples
Extreme-value moral hazard (principal–agent).